OpenStudy (anonymous):

The ratio of girls to boys in the classes held at th Fremont Dance Studio is 5 to 3. If 160 students take classes at the studio, how many are boys? a. 30 b. 60 c. 90 d. 100

OpenStudy (anonymous):

For ratio problems, please remember this rule. If the ratio of girls to boys is 5:3, then the number of girls is 5/8 and boys is 3/8. (We add 5 and 3 to get 8.) To find the number of boys, we multiply the total students by the fraction of boys, which is 160 x 3/8 = 60

OpenStudy (anonymous):

\[\frac{boys}{girls}=\frac{3}{5}\implies b=\frac{3}{8}160=60\]

OpenStudy (anonymous):

ohhh....i didn't even think to add 5 and 3. Thanks so much for your help!